# 6D-Diff: A Keypoint Diffusion Framework for 6D Object Pose Estimation

### 02 Jan 2024 v2

@@ 3.3 Training and Testing
In the first stage, to initialize the distribution $D_{K}$, we optimize the keypoints distribution initializer. Specifically, for each training sample, given the pre-selected $N$ 3D keypoints, we can obtain the ground-truth coordinates of the corresponding $N$ 2D keypoints using the ground-truth 6D object pose. Then for each keypoint, based on the corresponding ground-truth coordinates, we generate a ground-truth heatmap following [40] for training the initializer. Thus for each training sample, we generate $N$ ground-truth heatmaps. In this way, the loss function $L_{init}$ for optimizing the initializer can be formulated as:
$L_{init}={\Big{\|}\textbf{H}_{pred}-\textbf{H}_{GT}\Big{\|}}^{2}_{2}$ (7)
where $\textbf{H}_{pred}$ and $\textbf{H}_{GT}$ denote the predicted heatmaps and ground-truth heatmaps, respectively.
In the second stage, we optimize the diffusion model $M_{diff}$. For each training sample, to optimize $M_{diff}$, we perform the following steps. (1) We first send the input image into an off-the-shelf object detector [55] and then feed the detected ROI into the trained initializer to obtain $N$ heatmaps. Meanwhile, we can also obtain $f_{app}$. (2) We use the $N$ predicted heatmaps to initialize $D_{K}$, and leverage the EM-type algorithm to characterize $D_{K}$ as a MoC distribution $\hat{D}_{K}$. (3) Based on $\hat{D}_{K}$, we use the ground-truth keypoints coordinates $d_{0}$ to directly generate $M$ sets of ($\hat{d}_{1},...,\hat{d}_{K}$) (i.e., $\{\hat{d}^{i}_{1},...,\hat{d}^{i}_{K}\}^{M}_{i=1}$) via the forward process (Eq. 5). (4) Then, we aim to optimize the diffusion model $M_{diff}$ to recover $\hat{d}^{i}_{k-1}$ from $\hat{d}^{i}_{k}$ iteratively. Following previous diffusion works [17, 50], we formulate the loss $L_{diff}$ for optimizing $M_{diff}$ as follows ($\hat{d}^{i}_{0}=d_{0}$ for all $i$):
$L_{diff}=\sum^{M}_{i=1}\sum^{K}_{k=1}{\Big{\|}M_{diff}(\hat{d}^{i}_{k},f_{app},f^{k}_{D})-\hat{d}^{i}_{k-1}\Big{\|}}^{2}_{2}$ (8)
Testing. During testing, for each testing sample, by feeding the input image to the object detector and the keypoints distribution initializer consecutively, we can initialize $D_{K}$ and meanwhile obtain $f_{app}$. Then, we perform the reverse process. During the reverse process, we sample $M$ sets of noisy keypoints coordinates from $D_{K}$ (i.e., $\{d^{i}_{K}\}^{M}_{i=1}$) and feed them into the trained diffusion model. Here we sample $M$ sets of keypoints coordinates, because we are converting from a distribution ($D_{K}$) towards another distribution ($D_{0}$). Then the model iteratively performs the reverse steps. After $K$ reverse diffusion steps, we obtain $M$ sets of predicted keypoints coordinates (i.e., $\{d^{i}_{0}\}^{M}_{i=1}$). To obtain the final keypoints coordinates prediction $d_{0}$, we compute the mean of the $M$ predictions. Finally, we can solve 6D object pose using a PnP solver, like [42, 51].
@@ 3.4 Model Architecture
Our framework mainly consists of the diffusion model ($M_{diff}$) and the keypoints distribution initializer.
Diffusion Model $M_{diff}$. As illustrated in Fig. 3, our proposed diffusion model $M_{diff}$ mainly consists of a transformer encoder-decoder architecture. The object features $f_{app}$ are sent into the encoder for extracting context information to aid the reverse process in the decoder. $f^{k}_{D}$ and $\{d^{i}_{k}\}^{M}_{i=1}$ (or $\{\hat{d}^{i}_{k}\}^{M}_{i=1}$ during training) are sent into the decoder for the reverse process. Both the encoder and the decoder contain a stack of three transformer layers.
@@
More specifically, as for the encoder part, we first map $f_{app}\in\mathbb{R}^{16\times 16\times 512}$ through a 1 × 1 convolution layer to a latent embedding $e_{app}\in\mathbb{R}^{16\times 16\times 128}$. To retain the spatial information, following [57], we further incorporate positional encodings into $e_{app}$. Afterwards, we flatten $e_{app}$ into a feature sequence ($\mathbb{R}^{256\times 128}$), and send into the encoder. The encoder output $f_{enc}$ containing the extracted object information will be sent into the decoder to aid the reverse process. Note that during testing, for each sample, we only need to conduct the above computation process once to obtain the corresponding $f_{enc}$.
The decoder part iteratively performs the reverse process. For notation simplicity, below we describe the reverse process for a single sample $d_{k}$ instead of the $M$ samples ($\{{d}^{i}_{1},...,{d}^{i}_{K}\}^{M}_{i=1}$). Specifically, at the $k$-th reverse step, to inject the current step number ($k$) information into the decoder, we first generate the step embedding $f^{k}_{D}\in\mathbb{R}^{1\times 128}$ using the sinusoidal function following [17, 50]. Meanwhile, we use an FC layer to map the input $d_{k}\in\mathbb{R}^{N\times 2}$ to a latent embedding $e_{k}\in\mathbb{R}^{N\times 128}$. Then we concatenate $f^{k}_{D}$ and $e_{k}$ along the first dimension, and send into the decoder. By interacting with the encoder output $f_{enc}$ (extracted object information) via cross-attention at each layer, the decoder produces $f_{dec}$, which is further mapped into the keypoints coordinates prediction $d_{k-1}\in\mathbb{R}^{N\times 2}$ via an FC layer. Then we send $d_{k-1}$ back to the decoder as the input to perform the next reverse step.
Keypoints Distribution Initializer. The initializer adopts a ResNet-34 backbone, which is commonly used in 6D pose estimation methods [59, 51, 4]. To generate heatmaps to initialize the distribution $D_{K}$, we add two deconvolution layers followed by a 1 × 1 convolution layer after the ResNet-34 backbone, and then we obtain predicted heatmaps $\textbf{H}_{pred}\in\mathbb{R}^{N\times\frac{H}{4}\times\frac{W}{4}}$ where $H$ and $W$ denote the height and width of the input ROI image respectively. Moreover, the features outputted by the ResNet-34 backbone, combined with useful features obtained from other methods [51, 33], are used as the object features $f_{app}$.
@@ 4.1 Datasets & Evaluation Metrics
LM-O Dataset. The Linemod Occlusion (LM-O) dataset contains 1214 images and is a challenging subset of the LM dataset. In this dataset, around 8 objects are annotated on each image and the objects are often heavily occluded. Following [51, 4], we use both the real images from the LM dataset and the publicly available physically-based rendering (pbr) images [6] as the training images for LM-O. Following [59, 51], on LM-O dataset, we evaluate model performance using the commonly-used ADD(-S) metric. For this metric, we compute the mean distance between the model points transformed using the predicted pose and the same model points transformed using the ground-truth pose. For symmetric objects, following [61], the mean distance is computed based on the closest point distance. If the mean distance is less than 10% of the model diameter, the predicted pose is regarded as correct.
YCB-V Dataset. The YCB-V dataset is a large-scale dataset containing 21 objects and over 100k real images. The samples in this dataset often exhibit occlusions and cluttered backgrounds. Following [51, 4], we use both the real images from the training set of the YCB-V dataset and the publicly available pbr images as the training images for YCB-V. Following [59, 51], we evaluate model performance using the following metrics: ADD(-S), AUC (Area Under the Curve) of ADD-S, and AUC of ADD(-S). For calculating AUC, we set the maximum distance threshold to 10 cm following [61].
@@ 4.2 Implementation Details
We conduct our experiments on Nvidia V100 GPU. We set the number of pre-selected 3D keypoints $N$ to 128. During training, following [51, 32], we utilize the dynamic zoom-in strategy to produce augmented ROI images. During testing, we use the detected bounding box with Faster RCNN [47] and FCOS [55] provided by CDPNv2 [32]. The cropped ROI image is resized to the shape of $3\times 256\times 256$ ($H=W=256$). We characterize $D_{K}$ via a MoC model with 9 Cauchy kernels ($U=9$) for the forward diffusion process. We optimize the diffusion model $M_{diff}$ for 1500 epochs using the Adam optimizer [26] with an initial learning rate of 4e-5. Moreover, we set the number of sampled sets $M$ to 5, and the number of diffusion steps $K$ to 100. Following [51], we use Progressive-X [1] as the PnP solver. Note that during testing, instead of performing the reverse process with all the $K$ steps, we accelerate the process with DDIM [50], a recently proposed diffusion acceleration method. With DDIM acceleration, we only need to perform 10 steps to finish the reverse process during testing.
@@ 4.4 Ablation Studies
Figure 5: Qualitative results. Green bounding boxes represent the ground-truth poses and blue bounding boxes represent the predicted poses of our method. As shown, even facing severe occlusions, clutter in the scene or varying environment, our framework can still accurately recover the object poses, showing the effectiveness of our method for handling the noise and indeterminacy caused by various factors in object pose estimation. More qualitative results can refer to Supplementary.
We conduct extensive ablation experiments on the LM-O dataset. In these experiments, we report the model performance on ADD(-S) metric averaged over all the objects. More ablation studies can refer to Supplementary.
@@
Impact of denoising process. In our framework, we predict keypoints coordinates via performing the denoising process. To evaluate the efficacy of this process, we test three variants. In the first variant (Variant A), we remove the diffusion model $M_{diff}$ and predict keypoints coordinates directly from the heatmaps produced by the keypoints distribution initializer. The second variant (Variant B) has the same model architecture as our framework, but the diffusion model is optimized to directly predict the coordinates instead of learning the reverse process. Same as Variant B, the third variant (Variant C) is also optimized to directly predict coordinates without denoising process. For Variant C, we stack our diffusion model structure multiple times to produce a deep network, which has similar computation complexity with our framework. As shown in Tab. 3, compared to our framework, the performance of these variants significantly drops, showing that the effectiveness of our framework mainly lies in the designed denoising process.
Table 4: Evaluation on the effectiveness of the object features $f_{app}$.
@@
Impact of object features $f_{app}$. In our framework, we send the object features $f_{app}$ into the diffusion model $M_{diff}$ to aid the reverse process. To evaluate its effect, we test a variant in which we do not send $f_{app}$ into $M_{diff}$ (w/o $f_{app}$). As shown in Tab. 4, our framework performs better than this variant, showing that $f_{app}$ can aid $M_{diff}$ to get more accurate predictions.
@@
Impact of MoC design. During training, we model the distribution $D_{K}$ from the intermediate representation (heatmaps) as a MoC distribution $\hat{D}_{K}$, and train the diffusion model $M_{diff}$ to perform the reverse process from $\hat{D}_{K}$. To investigate the impact of this design, we evaluate two variants that train $M_{diff}$ in different ways. In the first variant (Standard diffusion w/o MoC), we train the model to start the reverse process from the standard Gaussian noise, i.e., following the basic forward process in Eq. 3 for model training. In the second variant (Heatmaps as condition), we still train the model to start denoising from the random Gaussian noise but we use the heatmaps as the condition for the reverse process. As shown in Tab. 5, our framework consistently outperforms both variants, showing effectiveness of the designed MoC-based forward process.
Runtime analysis. We test the runtime of our framework on a desktop with an AMD 3.90GHz CPU and an Nvidia 4090 GPU. The object detection with FCOS detector [55] takes 16.4 ms. The runtime of the keypoints distribution initializer is 20.3 ms. Then performing the reverse diffusion process takes 102.8 ms. Finally, for computing the object pose based on the prediction of the diffusion model, we use Progressive-X [1] as the PnP solver and it takes 58.2 ms. In this way, our framework totally needs around 197.7 ms to obtain the object pose.
@@ 5 Conclusion
In this paper, we have proposed a novel diffusion-based 6D object pose estimation framework, which effectively handles noise and indeterminacy in object pose estimation. In our framework, we formulate object keypoints detection as a carefully-designed reverse diffusion process. We design a novel MoC-based forward process to effectively utilize the distribution priors in intermediate representations. Our framework achieves superior performance on two commonly-used benchmarks.

### 22 Mar 2024 v3

@@ 3.3 Training and Testing
In the first stage, to initialize the distribution $D_{K}$, we optimize the keypoints distribution initializer. Specifically, for each training sample, given the pre-selected $N$ 3D keypoints, we can obtain the ground-truth coordinates of the corresponding $N$ 2D keypoints using the ground-truth 6D object pose. Then for each keypoint, based on the corresponding ground-truth coordinates, we generate a ground-truth heatmap following [42] for training the initializer. Thus for each training sample, we generate $N$ ground-truth heatmaps. In this way, the loss function $L_{\text{init}}$ for optimizing the initializer can be formulated as:
$L_{\text{init}}={\Big{\|}\textbf{H}_{\text{pred}}-\textbf{H}_{\text{GT}}\Big{\|}}^{2}_{2}$ (7)
where $\textbf{H}_{\text{pred}}$ and $\textbf{H}_{\text{GT}}$ denote the predicted heatmaps and ground-truth heatmaps, respectively.
In the second stage, we optimize the diffusion model $M_{\text{diff}}$. For each training sample, to optimize $M_{\text{diff}}$, we perform the following steps. (1) We first send the input image into an off-the-shelf object detector [57] and then feed the detected ROI into the trained initializer to obtain $N$ heatmaps. Meanwhile, we can also obtain $f_{\text{app}}$. (2) We use the $N$ predicted heatmaps to initialize $D_{K}$, and leverage the EM-type algorithm to characterize $D_{K}$ as a MoC distribution $\hat{D}_{K}$. (3) Based on $\hat{D}_{K}$, we use the ground-truth keypoints coordinates $d_{0}$ to directly generate $M$ sets of ($\hat{d}_{1},...,\hat{d}_{K}$) (i.e., $\{\hat{d}^{i}_{1},...,\hat{d}^{i}_{K}\}^{M}_{i=1}$) via the forward process (Eq. (5)). (4) Then, we aim to optimize the diffusion model $M_{\text{diff}}$ to recover $\hat{d}^{i}_{k-1}$ from $\hat{d}^{i}_{k}$ iteratively. Following previous diffusion works [18, 52], we formulate the loss $L_{\text{diff}}$ for optimizing $M_{\text{diff}}$ as follows ($\hat{d}^{i}_{0}=d_{0}$ for all $i$):
$L_{\text{diff}}=\sum^{M}_{i=1}\sum^{K}_{k=1}{\Big{\|}M_{\text{diff}}(\hat{d}^{i}_{k},f_{\text{app}},f^{k}_{D})-\hat{d}^{i}_{k-1}\Big{\|}}^{2}_{2}$ (8)
Testing. During testing, for each testing sample, by feeding the input image to the object detector and the keypoints distribution initializer consecutively, we can initialize $D_{K}$ and meanwhile obtain $f_{\text{app}}$. Then, we perform the reverse process. During the reverse process, we sample $M$ sets of noisy keypoints coordinates from $D_{K}$ (i.e., $\{d^{i}_{K}\}^{M}_{i=1}$) and feed them into the trained diffusion model. Here we sample $M$ sets of keypoints coordinates, because we are converting from a distribution ($D_{K}$) towards another distribution ($D_{0}$). Then the model iteratively performs the reverse steps. After $K$ reverse diffusion steps, we obtain $M$ sets of predicted keypoints coordinates (i.e., $\{d^{i}_{0}\}^{M}_{i=1}$). To obtain the final keypoints coordinates prediction $d_{0}$, we compute the mean of the $M$ predictions. Finally, we can solve for the 6D object pose using a PnP solver, like [44, 53].
@@ 3.4 Model Architecture
Our framework mainly consists of the diffusion model ($M_{\text{diff}}$) and the keypoints distribution initializer.
Diffusion Model $M_{\text{diff}}$. As illustrated in Fig. 3, our proposed diffusion model $M_{\text{diff}}$ mainly consists of a transformer encoder-decoder architecture. The appearance features $f_{\text{app}}$ are sent into the encoder for extracting context information to aid the reverse process in the decoder. $f^{k}_{D}$ and $\{d^{i}_{k}\}^{M}_{i=1}$ (or $\{\hat{d}^{i}_{k}\}^{M}_{i=1}$ during training) are sent into the decoder for the reverse process. Both the encoder and the decoder contain a stack of three transformer layers.
@@
More specifically, as for the encoder part, we first map $f_{\text{app}}\in\mathbb{R}^{16\times 16\times 512}$ through a 1 × 1 convolution layer to a latent embedding $e_{\text{app}}\in\mathbb{R}^{16\times 16\times 128}$. To retain the spatial information, following [59], we further incorporate positional encodings into $e_{\text{app}}$. Afterwards, we flatten $e_{\text{app}}$ into a feature sequence ($\mathbb{R}^{256\times 128}$), and send it into the encoder. The encoder output $f_{\text{enc}}$ containing the extracted object information will be sent into the decoder to aid the reverse process. Note that during testing, for each sample, we only need to conduct the above computation process once to obtain the corresponding $f_{\text{enc}}$.
The decoder part iteratively performs the reverse process. For notation simplicity, below we describe the reverse process for a single sample $d_{k}$ instead of the $M$ samples ($\{{d}^{i}_{1},...,{d}^{i}_{K}\}^{M}_{i=1}$). Specifically, at the $k$-th reverse step, to inject the current step number ($k$) information into the decoder, we first generate the step embedding $f^{k}_{D}\in\mathbb{R}^{1\times 128}$ using the sinusoidal function following [18, 52]. Meanwhile, we use an FC layer to map the input $d_{k}\in\mathbb{R}^{N\times 2}$ to a latent embedding $e_{k}\in\mathbb{R}^{N\times 128}$. Then we concatenate $f^{k}_{D}$ and $e_{k}$ along the first dimension, and send it into the decoder. By interacting with the encoder output $f_{\text{enc}}$ (extracted object information) via cross-attention at each layer, the decoder produces $f_{\text{dec}}$, which is further mapped into the keypoints coordinates prediction $d_{k-1}\in\mathbb{R}^{N\times 2}$ via an FC layer. Then we send $d_{k-1}$ back to the decoder as the input to perform the next reverse step.
Keypoints Distribution Initializer. The initializer adopts a ResNet-34 backbone, which is commonly used in 6D pose estimation methods [61, 53, 4]. To generate heatmaps to initialize the distribution $D_{K}$, we add two deconvolution layers followed by a 1 × 1 convolution layer after the ResNet-34 backbone, and then we obtain predicted heatmaps $\textbf{H}_{\text{pred}}\in\mathbb{R}^{N\times\frac{H}{4}\times\frac{W}{4}}$ where $H$ and $W$ denote the height and width of the input ROI image respectively. Moreover, the features outputted by the ResNet-34 backbone, combined with features obtained from methods [53, 35], are used as the object features $f_{\text{app}}$.
@@ 4.1 Datasets & Evaluation Metrics
LM-O Dataset. The Linemod Occlusion (LM-O) dataset contains 1214 images and is a challenging subset of the LM dataset. In this dataset, around 8 objects are annotated on each image and the objects are often heavily occluded. Following [53, 4], we use both the real images from the LM dataset and the publicly available physically-based rendering (pbr) images [6] as the training images for LM-O. Following [61, 53], on LM-O dataset, we evaluate the model performance using the commonly-used ADD(-S) metric. For this metric, we compute the mean distance between the model points transformed using the predicted pose and the same model points transformed using the ground-truth pose. For symmetric objects, following [63], the mean distance is computed based on the closest point distance. If the mean distance is less than 10% of the model diameter, the predicted pose is regarded as correct.
YCB-V Dataset. The YCB-V dataset is a large-scale dataset containing 21 objects and over 100k real images. The samples in this dataset often exhibit occlusions and cluttered backgrounds. Following [53, 4], we use both the real images from the training set of the YCB-V dataset and the publicly available pbr images as the training images for YCB-V. Following [61, 53], we evaluate the model performance using the following metrics: ADD(-S), AUC (Area Under the Curve) of ADD-S, and AUC of ADD(-S). For calculating AUC, we set the maximum distance threshold to 10 cm following [63].
@@ 4.2 Implementation Details
We conduct our experiments on an Nvidia V100 GPU. We set the number of pre-selected 3D keypoints $N$ to 128. During training, following [53, 34], we utilize the dynamic zoom-in strategy to produce augmented ROI images. During testing, we use the detected bounding box with Faster RCNN [49] and FCOS [57] provided by CDPNv2 [34]. The cropped ROI image is resized to the shape of $3\times 256\times 256$ ($H=W=256$). We characterize $D_{K}$ via a MoC model with 9 Cauchy kernels ($U=9$) for the forward diffusion process. We optimize the diffusion model $M_{\text{diff}}$ for 1500 epochs using the Adam optimizer [28] with an initial learning rate of 4e-5. Moreover, we set the number of sampled sets $M$ to 5, and the number of diffusion steps $K$ to 100. Following [53], we use Progressive-X [1] as the PnP solver. Note that during testing, instead of performing the reverse process with all the $K$ steps, we accelerate the process with DDIM [52], a recently proposed diffusion acceleration method. With DDIM acceleration, we only need to perform 10 steps to finish the reverse process during testing.
@@ 4.4 Ablation Studies
Figure 5: Qualitative results. Green bounding boxes represent the ground-truth poses and blue bounding boxes represent the predicted poses of our method. As shown, even facing severe occlusions, clutter in the scene or varying environment, our framework can still accurately recover the object poses, showing the effectiveness of our method for handling the noise and indeterminacy caused by various factors in object pose estimation.
We conduct extensive ablation experiments on the LM-O dataset, and we report the model performance on ADD(-S) metric averaged over all the objects.
@@
Impact of denoising process. In our framework, we predict keypoints coordinates via performing the denoising process. To evaluate the efficacy of this process, we test three variants. In the first variant (Variant A), we remove the diffusion model $M_{\text{diff}}$ and predict keypoints coordinates directly from the heatmaps produced by the keypoints distribution initializer. The second variant (Variant B) has the same model architecture as our framework, but the diffusion model is optimized to directly predict the coordinates instead of learning the reverse process. Same as Variant B, the third variant (Variant C) is also optimized to directly predict coordinates without denoising process. For Variant C, we stack our diffusion model structure multiple times to produce a deep network, which has similar computation complexity with our framework. As shown in Tab. 3, compared to our framework, the performance of these variants significantly drops, showing that the effectiveness of our framework mainly lies in the designed denoising process.
Table 4: Evaluation on the effectiveness of the object appearance features $f_{\text{app}}$.
@@
[…]
@@
Impact of MoC design. During training, we model the distribution $D_{K}$ from the intermediate representation (heatmaps) as a MoC distribution $\hat{D}_{K}$, and train the diffusion model $M_{\text{diff}}$ to perform the reverse process from $\hat{D}_{K}$. To investigate the impact of this design, we evaluate two variants that train $M_{\text{diff}}$ in different ways. In the first variant (Standard diffusion w/o MoC), we train the model to start the reverse process from the standard Gaussian noise, i.e., following the basic forward process in Eq. (3) for model training. In the second variant (Heatmaps as condition), we still train the model to start denoising from the random Gaussian noise but we use the heatmaps as the condition for the reverse process. As shown in Tab. 5, our framework consistently outperforms both variants, showing effectiveness of the designed MoC-based forward process.
@@ 5 Conclusion
In this paper, we proposed a novel diffusion-based 6D object pose estimation framework, which effectively handles noise and indeterminacy in object pose estimation. In our framework, we formulate object keypoints detection as a carefully-designed reverse diffusion process. We design a novel MoC-based forward process to effectively utilize the distribution priors in intermediate representations. Our framework achieves superior performance.
Acknowledgement. This work was supported by the National Research Foundation Singapore under the AI Singapore Programme (Award Number: AISG-100E-2023-121).